# Quantum Gravity, Hydrodynamics and Emergent Cosmology: A Collection of Perspectives

### v1

@@ 1 Introduction
This vision of hydrodynamics on superspace is likely universal. TGFT already represents a synthesis of multiple QG approaches (lattice quantum gravity, tensor models, loop quantum gravity) but similar frameworks emerge from other independent QG methodologies, indicating a shared foundation. For example, a superspace hydrodynamics model arises in loop quantum cosmology (LQC) Banerjee:2011qu within the separate universe approach Bojowald:2012wi and appears in models using minisuperspace quantization of supergravity’s exceptional symmetries Kleinschmidt:2009hv . A related dynamic is seen in the path integral formalism for covariant quantum gravity Giddings:1988wv , where effective superspace hydrodynamics emerge in the continuum limit Ambjorn:2009fm . While these commonalities are impressive, one should bear in mind that each QG formalism may still lead to unique dynamics and thus to characteristic observable consequences.
@@
Establishing such an overarching framework should provide robust explanations for long-standing cosmological puzzles but should certainly also advance the ability to derive predictions for observable QG effects in cosmology. However, it is already well-known that directly probing the physical conditions in which significant semi-classical and QG phenomena occur remains notoriously challenging. This difficulty has, in part, driven the pursuit of analog gravity models in condensed matter systems, particularly quantum fluids. In such systems, perturbations of the fluids’ hydrodynamic configurations behave analogously to matter fields on a curved geometry defined by hydrodynamic variables and are thus described by quantum field theory (QFT) in curved spacetime. This concept has fueled a wide array of theoretical studies of analog gravitational phenomena, including exotic effects such as black hole evaporation and cosmological particle production, with some of these phenomena even reproduced experimentally in laboratory settings Barcelo:2005fc . However, this approach faces a fundamental limitation: the analogy fails once the dynamical properties of the emergent geometry are considered. Specifically, the hydrodynamic equations governing the emergent metric do not map onto gravitational equations. The analog system is a non-relativistic fluid in flat space with preferred reference frames, in contrast to the background and frame independence required by general relativity (GR). Consequently, traditional analog models are constrained to simulate gravitational systems only at the kinematical level. Nonetheless, the editors of this collection propose that this limitation may be overcome with a paradigm shift in perspective: The hydrodynamic/cosmology mappings described above provide examples that circumvent this issue, demonstrating that gravitational dynamics and the hydrodynamics of quantum fluids can indeed align, with their symmetries matching as well. Therefore, to render the proposed framework testable, one should not recreate the spacetime manifold itself in the lab but superspace—the space of field configurations.
@@
This perspective motivated the collection’s structure and the selection of its contributors grouped into four thematic units exploring critical aspects of the superspace hydrodynamics framework: (a) Hydrodynamics/cosmology correspondence and link to analog gravity (contributors: Jibril Ben Achour, Salvatore Butera, and Tobias Haas), (2) Phase transitions, continuum limits and emergence in QG (contributors: Renata Ferrero, Roberto Percacci, Antonio D. Pereira, Dario Benedetti, Sebastian Steinhaus, and Johannes Thürigen), (3) Relational physics in gravity and QG (contributors: Flaminia Giacomini, Kristina Giesel, Joshua Kirklin, and Philipp Höhn), and (4) Emergent cosmology from QG (contributors: Mairi Sakellariadou, Martin Bojowald, Luca Marchetti and Robert Brandenberger). The first unit focuses on the vision of cosmology as hydrodynamics on superspace and recent developments in the field of analog gravity with the goal of discussing the change of perspective just described. The second unit centers on the use of mean-field and RG techniques to study via coarse-graining the continuum limit and emergence of spacetime in different QG approaches. The third unit explores recent advances on the relational strategy in gravity and quantum gravity to construct background-independent and diffeomorphism invariant observables. Finally, the fourth unit sheds light on the intimate relation between cosmology and quantum gravity as well as emergence of cosmology from quantum gravity approaches. These four thematic units reflect the interdisciplinary dimension of this workshop. Beyond this, note that the invited contributions by proponents of different QG approaches such as asymptotic safety in quantum gravity, CDT, spin foam gravity, loop quantum gravity, TGFT and string theory allow to map the plurality of the QG landscape and to discuss if and how one may derive the proposed framework from them.
@@
It is important to note that the editors’ vision of hydrodynamics on superspace as a framework for QG and cosmology may potentially not be shared by the various contributors to this collection whose perspective can be gathered from their individual contributions. Correspondingly, beyond the view of the organizers, this collection naturally reports on recent progress in the field of quantum gravity and fosters an exchange of ideas therein, in general. Finally, this collection is concluded on in the final section.
@@ 2 Map between hydrodynamics and cosmology
In this first part of the collection, we shall discuss several new results which suggest not only that the quantum cosmological dynamics can be reformulated as a hydrodynamical flow on the space of field configurations, also called super-space (or its extended version dubbed the lift hereafter) but also that this shall open the door to develop experimentally testable analog models for quantum cosmology.
This vision builds on i) mathematical results in symmetry-reduced general relativity and ii) on a paradigm shift in analog gravity. For this, consider the field of analog gravity. The ability to reproduce semi-classical effects from quantum field theory on curved backgrounds has remained for a long time a source of debate, mostly driven by a misleading name, starting from the field itself. Analog gravity is not about gravity per se but concerns a quantum field theory on a curved background. This background does not have to be a fixed spacetime representing part of a solution of general relativity, or any other theory of gravity. The only crucial point is that it has to be curved. In particular, this structure can be any other relevant space that one uses to encode the dynamics of a given field theory.
@@
The quantum dynamics of the homogeneous sector of cosmological geometries can be precisely described in this way. This is the standard mini-superspace approach. The quantum cosmological dynamics are described by field evolving on a curved background which is not the spacetime but the configuration space. This space describes all the cosmological metrics (mapped to a two-dimensional space labelled by the scale factor and a matter field in general) and admits in general a non-trivial curvature. Quantum cosmology is then described as a free complex scalar field theory on this curved background. Therefore, if one reproduces this curved background in the lab using a suitable analog system, such as a Bose-Einstein condensate, one could in principle contemplate the dynamics of the wave function of the universe as described in standard quantum cosmology. Nevertheless, in order to build such a quantum universe in the lab, several key notions have to be properly understood and controlled, starting from the dictionary between the observables of the fluid system and the quantum cosmological observables which are by construction relational (e.g. expectation values of the volume w.r.t the reference scalar field for example, see also Section 4 for further information on the relationalist perspective), the translation of the meaning of the boundary conditions in the experimental in terms of cosmology, a clear understanding of the gauge symmetry in the cosmological set-up (choice of clock), etc..
@@
Nevertheless, if this picture is successfully realized and these various obstacles overcome, it will provide a fascinating perspective for both quantum gravity and analog gravity. On the one hand, we shall have access to experimentally controlled systems mimicking the very early universe, which can be tuned and reproduced at will. On the other hand, it will open the door to bypass a key limitation of analog gravity, namely that the curved background reproduced so far in any system does not contain any dynamical information on the system. By giving up the focus on reproducing the spacetime structure but instead focusing on auxiliary spaces such as the space of field configuration, analog gravity will emancipate itself from its current status, with potential new phenomena to be observed in the lab.
In the following, we discuss three interconnected pieces which advocate for this new point of view, carefully presenting the remaining gaps to be filled before achieving a consistent realization of the paradigm shift. The first contribution summarizes the recent mathematical results on symmetries of gravitational mini-superspaces which shall allow the construction of a dictionary between (quantum) fluids and quantum symmetry reduced gravitational systems, while the two other contributions report on recent progresses in analog gravity to i) take into account backreaction effects and ii) simulate a Friedman-Lemaître-Roberston-Walker geometry and the behavior of quantum fields on it in the lab. These recent progresses underline fundamental questions which are relevant to construct and progress towards the paradigm shift advocated in this manuscript.
@@ 2.1 Quantum cosmology as hydrodynamics in field space
This point of view resonates with different approaches of quantum gravity which aim at recovering the cosmological dynamics as an hydrodynamical approximation of a condensate state of pre-geometric building blocks. Important efforts have been developed in this direction in the group field theory (GFT) approach to quantum gravity where the bouncing cosmological dynamics with the right semi-classical limit was recovered following this scheme Gielen:2013naa ; Oriti:2016qtz ; Pithis:2019tvp ; Jercher:2021bie . Now, such derivation requires several layers of approximations which need to be controlled in order to fully develop a consistent emergent cosmology from these GFT models. While efforts are devoted to further refine and characterize this mean-field approximation Marchetti:2020xvf ; Marchetti:2022igl ; Marchetti:2022nrf ; Dekhil:2024ssa ; Dekhil:2024djp , it is interesting to see that one can extract directly from the IR regime symmetry-based-arguments which reinforce the picture of quantum cosmology as emerging from an hydrodynamical description in field space of a many-body system. See Oriti:2024qav for a recent account on this approach.
To conclude, it should be stressed that this result also suggests that the standard WdW dynamics used in quantum cosmology (but also in quantum black hole mechanics) could be generalized by non-linear interation of the wave function of the geometry just as the free Schrödinger equation can be extended to Gross-Pitaevskii equation when describing condensates. Indeed, the Schrödinger invariance allows one to uniquely select non-linear corrections which preserves this symmetry, giving rise to a one parameter family of non-linear WdW equations BenAchour:2023dgj . Therefore, it would be useful to explore what could be the phenomenological consequences of this new non-linear extension of the WdW dynamics in quantum cosmology. Nevertheless, more work is needed to understand how this physical symmetry can be contemplated and exploited without fixing the gauge. To that end, it would be useful to extend the Einseinhart-Duval lift to fully time-reparametrization mechanical systems, by understanding how the lapse can be included in this geometric framework.
@@ 2.3 A Universe in Heidelberg
Following this perspective’s spirit and given the deep connection between hydrodynamic and cosmological models in their symmetries BenAchour2022 , see also Section 2.1, one might consider the fascinating possibility of hydrodynamic analogies for quantum cosmology models. (See the first contribution for details). These models feature an equation of motion for the wave function of the universe, the Wheeler-DeWitt (WDW) equation Hartle1983 ; Kiefer2022 , which might be quantum simulated by the dynamics of acoustic excitations in ultracold atom experiments similar to Vierman2022 . In this way, predictions of quantum cosmology models could ultimately be rendered accessible to experimental tests Carney2019 and analog gravity experiments would become the window into one of the most concealed pieces of our world: the quantum nature of spacetime.
@@ 3 Phase transitions, continuum limits and emergence in quantum gravity
This section aims to present recent advances and highlight outstanding challenges within these approaches and establishes connections to Section 4 on relationalism and Section 5 on deriving effective cosmology from quantum gravity.

### v2

@@ 1 Introduction
This vision of hydrodynamics on superspace is likely universal. TGFT already represents a synthesis of multiple QG approaches (lattice quantum gravity, tensor models, loop quantum gravity) but similar frameworks emerge from other independent QG methodologies, indicating a shared foundation. For example, a superspace hydrodynamics model arises in loop quantum cosmology (LQC) Banerjee:2011qu within the separate universe approach Bojowald:2012wi and appears in models using mini-superspace quantization of supergravity’s exceptional symmetries Kleinschmidt:2009hv . A related dynamic is seen in the path integral formalism for covariant quantum gravity Giddings:1988wv , where effective superspace hydrodynamics emerge in the continuum limit Ambjorn:2009fm . While these commonalities are impressive, one should bear in mind that each QG formalism may still lead to unique dynamics, and thus to characteristic observable consequences.
@@
Establishing such an overarching framework should provide robust explanations for long-standing cosmological puzzles, but should certainly also advance the ability to derive predictions for observable QG effects in cosmology. However, it is already well known that directly probing the physical conditions in which significant semi-classical and QG phenomena occur remains notoriously challenging. This difficulty has, in part, driven the pursuit of analog gravity models in condensed matter systems, particularly quantum fluids. In such systems, perturbations of the fluids’ hydrodynamic configurations behave analogously to matter fields on a curved geometry defined by hydrodynamic variables and are thus described by quantum field theory (QFT) in curved spacetime. This concept has fueled a wide range of theoretical studies of analog gravitational phenomena, including exotic effects such as black hole evaporation and cosmological particle production, with some of these phenomena even reproduced experimentally in laboratory settings Barcelo:2005fc . However, this approach faces a fundamental limitation: the analogy fails once the dynamical properties of emergent geometry are considered. Specifically, the hydrodynamic equations governing the emergent metric do not map onto gravitational equations. The analog system is a non-relativistic fluid in flat space with preferred reference frames, in contrast to the background and frame independence required by general relativity (GR). Consequently, traditional analog models are constrained to simulate gravitational systems only at the kinematical level. Nonetheless, the editors of this collection propose that this limitation may be overcome with a paradigm shift in perspective: The hydrodynamic/cosmology mappings described above provide examples that circumvent this issue, demonstrating that gravitational dynamics and the hydrodynamics of quantum fluids can indeed align, with their symmetries matching as well. Therefore, to render the proposed framework testable, one should not recreate the spacetime manifold itself in the lab but superspace—the space of field configurations.
@@
This perspective motivated the collection’s structure and the selection of its contributors grouped into four thematic units exploring critical aspects of the superspace hydrodynamics framework: (a) Hydrodynamics/cosmology correspondence and link to analog gravity (contributors: Jibril Ben Achour, Salvatore Butera, and Tobias Haas), (2) Phase transitions, continuum limits, and emergence in QG (contributors: Renata Ferrero, Roberto Percacci, Antonio D. Pereira, Dario Benedetti, Sebastian Steinhaus, and Johannes Thürigen), (3) Relational physics in gravity and QG (contributors: Flaminia Giacomini, Kristina Giesel, Joshua Kirklin, and Philipp Höhn), and (4) Emergent cosmology from QG (contributors: Mairi Sakellariadou, Martin Bojowald, Luca Marchetti, and Robert Brandenberger). The first unit focuses on the vision of cosmology as hydrodynamics on superspace and recent developments in the field of analog gravity with the goal of discussing the change of perspective just described. The second unit centers on the use of mean-field and RG techniques to study via coarse-graining the continuum limit and the emergence of spacetime in different QG approaches. The third unit explores recent advances on the relational strategy in gravity and quantum gravity to construct background-independent and diffeomorphism-invariant observables. Finally, the fourth unit sheds light on the intimate relation between cosmology and quantum gravity, as well as emergence of cosmology from quantum gravity approaches. In particular, a better understanding of this cosmic emergence process goes naturally hand in hand with progress in condensed matter physics, renormalization, phase transitions and relational physics. This progress must not only be technical, but also conceptual, as QG forces us to revisit these aspects of classical and quantum physics from its own perspective, stripping them of all the redundant, spacetime based information. This offers a unique chance of cross-fertilization and interdisciplinarity between QG and other fields in physics, as shown in this collection.
However, it is important to emphasize that the contributions can be viewed as standalone perspectives on their respective topics. It was up to the individual contributors to explain whether and how they their research direction could more directly contribute to the proposed framework. Nevertheless, as outlined in the introductory remarks, they clearly inform the framework’s development through their results, as further elaborated in the sectional introductions. Beyond this, note that the invited contributions by proponents of different QG approaches such as asymptotic safety in quantum gravity, causal dynamical triangulations (CDT), spin foam gravity, loop quantum gravity, TGFT and string theory allow one to map the plurality of the QG landscape and to discuss if and how they relate to this framework.
@@
That being said, it important to note that the editors’ vision of hydrodynamics on superspace as a framework for QG and cosmology may potentially not be shared by the various contributors to this collection whose perspective can be gathered from their individual contributions. Correspondingly, beyond the view of the organizers, this collection naturally reports on recent progress in the field of quantum gravity and fosters an exchange of ideas therein, in general. Finally, this collection is concluded in the final section.
@@ 2 Map between hydrodynamics and cosmology
In this first part of the collection, we shall discuss several new results which suggest not only that the quantum cosmological dynamics can be reformulated as a hydrodynamical flow on the space of field configurations, also called superspace (or its extended version dubbed the lift hereafter), but also that this shall open the door to develop experimentally testable analog models for quantum cosmology.
This vision builds on i) mathematical results in symmetry-reduced general relativity and ii) on a paradigm shift in analog gravity. For this, consider the field of analog gravity at first. The ability to reproduce semi-classical effects from quantum field theory on curved backgrounds has remained for a long time a source of debate, mostly driven by a misleading name, starting from the field itself. Analog gravity is not about gravity per se but concerns a quantum field theory on a curved background. This background does not have to be a fixed spacetime representing part of a solution of general relativity, or any other theory of gravity. The only crucial point is that it has to be curved. In particular, this structure can be any other relevant space that one uses to encode the dynamics of a given field theory.
@@
The quantum dynamics of the homogeneous sector of cosmological geometries can be precisely described in this way. This is the standard mini-superspace approach. The quantum cosmological dynamics are described by field evolving on a curved background which is not the spacetime but the configuration space. This space describes all the cosmological metrics (mapped to a two-dimensional space labelled by the scale factor and a matter field in general) and admits in general a non-trivial curvature. Quantum cosmology is then described as a free complex scalar field theory on this curved background. Therefore, if one reproduces this curved background in the lab using a suitable analog system, such as a Bose-Einstein condensate, one could in principle contemplate the dynamics of the wave function of the universe as described in standard quantum cosmology. Nevertheless, in order to build such a quantum universe in the lab, several key notions have to be properly understood and controlled, starting from the dictionary between the observables of the fluid system and the quantum cosmological observables which are by construction relational (e.g. expectation values of the volume w.r.t the reference scalar field for example, see also Section 4 for further information on the relationalist perspective and Section 5.3 for an explicit discussion of this matter in the context of the TGFT condensate cosmology approach), the translation of the meaning of the boundary conditions in the experimental in terms of cosmology, a clear understanding of the gauge symmetry in the cosmological set-up (choice of clock), etc..
@@
Nevertheless, if this picture is successfully realized and these various obstacles overcome, it will provide a fascinating perspective for both quantum gravity and analog gravity. On the one hand, we shall have access to experimentally controlled systems mimicking the very early universe, which can be tuned and reproduced at will. On the other hand, it will open the door to bypass a key limitation of analog gravity, namely that the curved background reproduced so far in any system does not contain any dynamical information on the system. By giving up the focus on reproducing the spacetime structure but instead focusing on auxiliary spaces such as the space of field configurations, analog gravity will emancipate itself from its current status to be constrained to simulate gravitational systems only at the kinematical level, with potential new phenomena to be observed in the lab.
In the following, we discuss three interconnected pieces which advocate for this new point of view, carefully presenting the remaining gaps to be filled before achieving a consistent realization of the paradigm shift. The first contribution, Section 2.1, summarizes the various recent results regarding the shared symmetries between hydrodynamics and symmetry-reduced gravitational superspaces, relevant both for cosmology but also for black holes mechanics. This shall facilitate the construction of a dictionary between (quantum) fluids and quantum symmetry reduced gravitational systems. This dictionary suggests that the background configuration of the hydrodynamic flow should be understood as representing not a given spacetime geometry, but an auxiliary space encoding the space of field configurations and, in particular, the dynamical symmetries of the gravitational system. This key point opens the door for analog models to emancipate themselves from their gravity-oriented interpretation and to encode dynamical information on the system of interest at the level of the background. This idea naturally leads to interpreting (quantum) cosmology as a hydrodynamical flow on the space of field configurations. The second contribution, Section 2.2, reports on recent progress in analog gravity to take into account backreaction effects of quantum fluctuations onto the background geometry. In a sense, this effort goes in the same direction as it aims at encoding effects of the dynamics on the background geometry, unfreezing it from its current kinematical status. Finally, the third contribution, Section 2.3, reviews recent developments in analog gravity to set up suitable analog models tailored to cosmological settings, in particular to simulate a Friedmann-Lemaître-Roberston-Walker geometry and the behavior of quantum fields on it in the lab. Preliminary results aiming at providing an analog BEC model for quantum cosmology and reproducing the WDW dynamics have shown that the geometry of the field space takes a form very close to a cosmological spacetime geometry. It could very well be the case that the current efforts devoted to realizing an analog model for expanding spacetime geometries find another use when constructing such analog models for quantum cosmology. [^2] For work complementing the content of Sections 2.2 and 2.3, we refer the reader to Refs.  Schutzhold:2005ex ; Fischer:2005iy ; Baak:2022hum ; Pal:2024qno and Tian:2017cda ; Ribeiro:2022gln ; Cha:2016esj , respectively. These recent progresses underline fundamental questions which are relevant to construct and progress towards the paradigm shift advocated in this manuscript.
@@ 2.1 Quantum cosmology as hydrodynamics in field space
This point of view resonates with different approaches of quantum gravity which aim at recovering the cosmological dynamics as an hydrodynamical approximation through coarse-graining of pre-geometric building blocks. Important efforts have been developed in this direction in the group field theory (GFT) approach to quantum gravity where the bouncing cosmological dynamics with the right semi-classical limit was recovered following this scheme Gielen:2013naa ; Oriti:2016qtz ; Pithis:2019tvp ; Jercher:2021bie , see in particular Section 5.3 for a lightning review of the TGFT condensate cosmology approach. Now, such derivation requires several layers of approximations which need to be controlled in order to fully develop a consistent emergent cosmology from these GFT models. While efforts are devoted to further refine and characterize this coarse-grained perspective corresponding to a mean-field approximation Marchetti:2020xvf ; Marchetti:2022igl ; Marchetti:2022nrf ; Dekhil:2024ssa ; Dekhil:2024djp , it is interesting to see that one can extract directly from the IR regime symmetry-based-arguments which reinforce the picture of quantum cosmology as emerging from an hydrodynamical description in field space of a many-body system. These developments demonstrate that TGFT condensate cosmology is a specific instantiation of the hydrodynamics on superspace framework, featured in the general introduction to this collection. We refer the reader to Ref. Oriti:2024qav for a detailed presentation of this new framework.
To conclude, it should be stressed that this result also suggests that the standard WDW dynamics used in quantum cosmology (but also in quantum black hole mechanics) could be generalized by non-linear interaction of the wave function of the geometry just as the free Schrödinger equation can be extended to Gross-Pitaevskii equation when describing condensates. Indeed, the Schrödinger invariance allows one to uniquely select non-linear corrections which preserves this symmetry, giving rise to a one parameter family of non-linear WDW equations BenAchour:2023dgj . Therefore, it would be useful to explore what could be the phenomenological consequences of this new non-linear extension of the WDW dynamics in quantum cosmology. Nevertheless, more work is needed to understand how this physical symmetry can be contemplated and exploited without fixing the gauge. To that end, it would be useful to extend the Einseinhart-Duval lift to fully time-reparametrization mechanical systems, by understanding how the lapse can be included in this geometric framework.
@@ 2.3 A Universe in Heidelberg
Following this perspective’s spirit of ”hydrodynamics on superspace” and given the deep connection between hydrodynamic and cosmological models in their symmetries BenAchour2022 , see also Section 2.1, one might consider the fascinating possibility of hydrodynamic analogies for quantum cosmology models, see the first contribution for details. The equation of motion for the wave function of the universe, the Wheeler-DeWitt (WDW) equation Hartle1983 ; Kiefer2022 , which might be quantum simulated by the dynamics of acoustic excitations in ultracold atom experiments similar to Vierman2022 - when viewed through the lens of the hitherto discussed superspace theory. In this way, predictions of quantum cosmology models could ultimately be rendered accessible to experimental tests Carney2019 and analog gravity experiments would become the window into one of the most concealed pieces of our world: the quantum nature of spacetime.
@@ 3 Phase transitions, continuum limits and emergence in quantum gravity
The goal of this section is to present recent advances and to highlight outstanding challenges within these approaches. In particular, each contribution can be seen as a compact standalone introduction to its respective topic. Next to this, it also establishes connections to Section 4 on relationalism and Section 5 on deriving effective cosmology from quantum gravity. In view of the hydrodynamics on superspace framework advertised by the editors, this section discusses in part the topic of emergence in quantum gravity from discrete structures. In particular, Sections 3.1 and 3.3 advertise the use of mean-field techniques in the CDT and TGFT approaches to this end. While it is left to future investigations, how the results presented in Section 3.1 could fit in detail into this novel framework, the results on phase transitions and condensate states in TGFT, displayed by Section 3.3, directly support this approach as an instantiation of it. This is because the mean-field hydrodynamics of such states can be mapped to an effective continuum cosmological dynamics Gielen:2013kla ; Oriti:2016qtz ; Pithis:2019tvp ; Jercher:2021bie , see especially Section 5.3 for an elaboration of this matter. Also, the cosmological dynamics of effective spin foam models resulting from coarse-graining, see Section 3.2, may potentially fit into this framework. Clearly, the perspective discussed in Section 3.4 on asymptotic safety approach is different in this regard since it is based on fields living on continuum spacetime. However, the investigation of the phase structure therein and in particular the exploration of the continuum limit is a shared challenge among all of these approaches and shared tools, like the renormalization group, are successfully employed to this end. Beyond this, in anticipation of Section 4 on the relationalist perspective, Section 3.4 reports on recent efforts to employ relational frames in the asymptotic safety approach which is a common tool in quantum gravity approaches to construct gauge-invariant observables and to deparametrize the full dynamics. This method plays a highly important role in the hydrodynamics on superspace framework where spacetime is expected to be recovered as an emergent and approximate structure using it.